Pearl Inc. issues 500 shares of $10 par value common stock and 100 shares of $100 par value preferred stock for a lump sum of $121,000.

(a) Prepare the journal entry for the issuance when the market price of the common shares is $164 each and market price of the preferred is $205 each.
(b) Prepare the journal entry for the issuance when only the market price of the common stock is known and it is $212 per share.


(Round answers to 0 decimal places, e.g. $1,225. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Account Titles and Explanation

Debit

Credit

(a)

enter an account title for case A enter a debit amount enter a credit amount
enter an account title for case A enter a debit amount enter a credit amount
enter an account title for case A enter a debit amount enter a credit amount
enter an account title for case A enter a debit amount enter a credit amount
enter an account title for case A enter a debit amount enter a credit amount

(b)

enter an account title for case B enter a debit amount enter a credit amount
enter an account title for case B enter a debit amount enter a credit amount
enter an account title for case B enter a debit amount enter a credit amount
enter an account title for case B enter a debit amount enter a credit amount
enter an account title for case B enter a debit amount enter a credit amount
Answer #1

Answer a.

No. Account Titles and Explanation Debit Credit
(a) Cash 121,000
Common Stock (500 shares * $10) 5,000
Paid-in Capital in Excess of Par - Common Stock (96,800 - 5,000) 91,800
Preferred Stock (100 shares * $100) 10,000
Paid-in Capital in Excess of Par - Preferred Stock (24,200 - 10,000) 14,200
(To record issuance at a lump sum)

Working Note:

Market value of common stock issued (500 shares * $164) 82,000
Market value of preferred stock issued (100 shares * $205) 20,500
Total $102,500

Allocation to common stock = (82,000 / 102,500) * 121,000 = $96,800

Allocation to common stock = (20,500 / 102,500) * 121,000 = $24,200

Answer b.

No. Account Titles and Explanation Debit Credit
Cash 121,000
Common Stock (500 shares * $10) 5,000
Paid-in Capital in Excess of Par - Common Stock (106,000 - 5,000) 101,000
Preferred Stock (100 shares * $100) 10,000
Paid-in Capital in Excess of Par - Preferred Stock (15,000 - 10,000) 5,000
(To record issuance at a lump sum)

Working Note:

Lump sum amount 121,000
Less : Allocated to common stock (500 shares * $212) 106,000
= Allocated to preferred stock 15,000